# Precomputed Dominant Resource Fairness

### v3

@@ 1 Introduction
Dominant Resource Fairness (DRF) is a resource allocation scheme, which aims to provide fairness of allocation among users with heterogeneous demands, in a time-sharing system with multiple resource types, such as computing clouds [1]. The services offered by these systems may include processing power, memory, and bandwidth, all of which are expected to be distributed fairly among the clients, according to some definition of fairness.
Although the problem has been well-studied for single resource types [2, 3], prior to DRF, less attention was paid to the multi-resource scenarios. The general approach was to adopt well established solutions for single resource allocation, such as Max-min Fairness (MF) and its derivatives, for a fixed bundle with differing amounts of resources from each resource type; e.g. standardised virtual machines. But this approach leads to inefficiencies, since the needs of different users are heterogeneous, where one needs predominantly one resource type, the other needs another. In the end, users ended up with unused resources that come in the bundle with the resources they need, where at the same time there were other users in the system in demand of those unused resources.
@@
In the present paper, we analyse the DRF algorithm as introduced in [1] in closer view, and based on the observations gathered thereby, introduce a new algorithm to approximate the DRF allocation in fewer steps. We call this new algorithm Precomputed Dominant Resource Fairness (PDRF), after its main working principle of precomputing the main allocation loop of DRF, and assigning resources to users without going through the tedious iterations of the loop.
One study that also needs to be mentioned at the onset, is [4], the method and the findings of which closely resemble to that of the present study, with some key differences that allow us to implement our newly proposed algorithm. The analysis method and the algorithm they propose, the Extended Dominant Resource Fairness (EDRF) are similarly based on precomputing the final allocation of DRF, but they modeled the underlying linear programming problem, rather than the algorithm itself. They construct a theoretical framework that allows them to provide more rigorous proofs of DRF properties, which is the main objective of the article. We elaborate on this in Section 3.2.
@@ 2 Related Work
DRF is first proposed in [1], where they introduce the new notion of establishing an MF allocation among dominant shares, rather than a fixed resource bundle of different resource types. They also provide an algorithm to carry out the distribution, and produce empirical data to demonstrate the performance of the algorithm. In their theoretical sections, they compare the virtues of the newly introduced allocation scheme to the ones existing in the economy literature, for this is as much a problem in the field of economy as it is in the field of computer science. Their results clearly demonstrate that DRF possesses various advantages over its alternatives.
The model is challenged in [5]. They developed an alternative fairness notion, which they refer to as no justified complaints, based on the assumption that the fairness of distribution of resources should be assessed from a global system point of view, considering the bottleneck property of resources. A resource is a bottleneck resource, if the availability of this resource is limited and there is a race condition for accessing it. They developed the Bottleneck Based Fairness (BBF) algorithm to implement this paradigm as an alternative to DRF, with relative advantages of each algorithm in different use cases.
Not only challenged, but DRF has also been adopted, generalised, contextualised, or extended to different cases. For example, Gutman and Nisan [6], introduced an economical framework to compare and contrast the two systems described above. While doing so they also critically assessed both approaches from an economic point of view. Joe-Wong and colleagues [7] investigates the trade-off between fairness and efficiency, providing a unifying framework with two families of fairness functions, DRF being a special case in one of them. Kash [8] takes a scenario in which the complete information on demands is not available, since users arrive and leave the system, and develops a Dynamic DRF (DDRF) to address the problem. The DDRF algorithm is improved for its performance in [9].
Finally, DRF has been reformulated by Parkes and colleagues [4] to EDRF, in order to study its properties more rigorously. EDRF is also a solution to a special case, where the resources, as well as user demands are well-divisible, i.e. fractions of demands can be allocated to users, but for obvious reasons this is not widely applicable to real life scenarios, where half a task does not provide any added utility to the users. In the same article, they also introduce Sequential Min-max Algorithm to address the indivisible demands scenario. One more contribution of EDRF is that it addresses the situations where there are users who submit zero demands for one or more resources.
@@ 3.1 Dominant Resource Fairness
The DRF allocation scheme is build on the concept of dominant shares. The dominant share for each user is defined as the maximum of the fraction of her demands for different resource types to their respective reserves in the system. As exemplified in [1], in a system with $9$ CPUs and $18$ GB memory, the dominant share, $ds$ for user A, with the demand vector $\left\langle 1,4\right\rangle$ is memory, since the fractional demands, $fd$, of the user are $1/9$ and $4/18=2/9$, respectively. Similarly, the $ds$ of a user with the demand vector $\left\langle 3,1\right\rangle$ is for CPU, since $fd$’s are $3/9$ and $1/18$, respectively. The question then is, what is a fair allocation between the users, whose demands for the resources are defined as such.
If the number of tasks allocated to users A and B is denoted by $x$ and $y$, respectively, the DRF allocation is defined as the solution to the optimising problem below [1, p. 4]:
@@
As seen in the example above, DRF adopts progressive filling strategy, by aiming to keep the growth of allocation among different agents at the same pace. DRF reduces to MF, and also to BBF [5], in cases where all users’ dominant resources are the same.
@@ 3.2 Extended Dominant Resource Fairness
The study by [4] introduces an alternative view of DRF. Instead of an iterative algorithm, they aim to precompute the resulting distribution, with a linear program. For doing so, they employ two notions in the setting. First, in a given user demand vector, the ratio between the entries for different resources is fixed. Second, the operating principles of progressive filling algorithm dictates that the pace at which the allocation for dominant shares of different users grow should be equated.
@@ 4 Cyclic Structure of the DRF Main Loop
In a system of $n$ users, the length of the cycle and the number of occurrences for each individual demand are given by
@@
where $\lvert l\rvert$ denotes the length of the cycle, $\operatorname{lcm}DS$ denotes the least common multiple of the set of all dominant shares, $DS$, and $ds_{i}$ is the dominant share of user $i$. The individual occurrences of each demand within the cycle is given by the $i$-thth term of the summation for user $i$. In the cases where all other dominant shares are integer fractions of $ds^{*}$, like in the example above, since $lcmS$ is equal to $ds^{*}$, the latter part is equivalent to the ratio of each $ds_{i}$ to $ds^{*}$, and the length of the cycle is equal to the sum of all these ratios, plus $1$, for the single occurrence of $ds^{*}$.
The ratio of lower demands to the $ds^{*}$ is still useful, in cases where integer fraction assumption does not hold. In those cases, further subcycles emerge, characterised by the single occurrence of $ds^{*}$, followed by a differing number of all other demands. The most frequently occurring subcycle is characterised by a single occurrence of the demand with $ds^{*}$, followed by all other demands occurring for times equal to the integer part of the ratio of their dominant share to $ds^{*}$. That is to say:
$\lvert l^{*}\rvert=\sum_{i=1}^{n}\left\lfloor\frac{ds^{*}}{ds_{i}}\right\rfloor$
where $\lvert l^{*}\rvert$ denotes the length of, what we would like to call, the basic subcycle. The other subcycles are characterised by including an additional occurrence of one or more of the lower dominant share demands, the frequency of which is determined by the decimal part of the ratio of $ds^{*}$ to each dominant share, i.e. the $i$-thth term of the expression above for demand $i$. Trivially, the first subcycle will always have an extra occurrence, if the ratio has a decimal part.
Whenever a multiple of the decimal part reaches or exceeds an integer number, another allocation for the given demand shows up in the corresponding subcycle. For example, in a system with two demands, whose dominant share ratio is $1.25$, the basic subcycle will consist of each demand taking $1$ allocation, and every $4th$ subcycle will include one more allocation for the lower dominant share, making it $1$ for the highest dominant share and $2$ for the lower. In cases where the decimal summations exceed an integer, such as in the case of a ratio of $1.4$, there will be $2$ different periods of extra occurrence; in this case, after the first subcycle, every $2nd$ and then $3rd$ subcycle iterations.
Systems with more than two users, with all pairwise dominant share ratios having decimal parts, and not necessarily as nice as the ones presented in the examples above, lead to a cornucopia of subcycles, with differing and overlapping extra occurrence periods for each non-highest dominant share demand. Luckily, we will not need to identify all of these for the purposes pursued in the subsequent sections.
@@ 5 Precomputed Dominant Resource Fairness
The pseudocode of PDRF is presented in Algorithm 1. A sample implementation can be accessed at this github repository [10].
@@ 6 Results
$1000$ users are assigned random demands, and resource reserves are likewise assigned randomly. We tested for various numbers of resource types, demand intervals and resource reserve intervals. Here, we are presenting a sample which we consider representative. We run each test $1000$ times, and collected the averages, standard deviations, and maximums, presented in Tables 1 - 4. The scripts that we used to run the tests can be accessed at the same repository, here [10].
@@ 9.2 Pareto Efficiency
The proof is based on the claim that ”progressive filling for DRF is equivalent to the algorithm presented in Figure 11111Since Figure $1$ is a scatter plot, we assume what is meant is Algorithm $1$. [1, p. 13]. Although it is indicated in the preceding paragraph that the progressive filling algorithm eliminates the saturated demands, no removal condition is given in Algorithm $1$, and neither it is mentioned in the sections and examples describing the algorithm. This might be a minor point eluding the authors’ attention or it might have been assumed that the selection process automatically excludes those demands, but the resulting algorithm as such, cannot achieve Pareto Efficiency in, though rare might they be, certain demand distributions and resource reserves.
@@
An example we rather stumbled upon, which led to this analysis, consists of the same demand vectors in the original article, user A with the demand vector $\left\langle 1,4\right\rangle$ and user B with $\left\langle 3,1\right\rangle$, and the available capacities $\left\langle 59,19\right\rangle$. In such a setting, after 10 iterations, the algorithm halts allocating $\left\langle 2,8\right\rangle$ and $\left\langle 24,8\right\rangle$, respectively, and $\left\langle 33,3\right\rangle$ resources still available, which is sufficient for $3$ more allocations for user B, but since the least allocated dominant share belongs to A, and his demand is saturated, the algorithm cannot proceed. A short simulator script with this example can be found at the same github repository [10].
@@ 9.3 Weighting Policies
A weighting policy, assigns weights to users, and allocates the resources in proportion to these weights. Fairness, in turn, is defined over the allocations scaled by these weights. In [1], the question of weighting is addressed briefly, pointing to the fact that it suffices for such a generalisation to assign a weight vector to each user, $W_{i}=\left\langle w_{i1},w_{i2},w_{i3},\dots,w_{im}\right\rangle$, and compute the dominant shares with $ds_{i}=\max_{j}\left\{u_{ij}/w_{ij}\right\}$. In [4], weights are further defined to be normalised, such that $\sum_{i}w_{ir}=1$, for each resource $r$, and they use another $\rho$” variable, at the construction of which the weights scale the normalised demands, to be used in their calculations. Again, it is notationwise involved and will not be elaborated on in more detail for the sake of simplicity.

### v4

@@ 1 Introduction
Dominant Resource Fairness (DRF) is a resource allocation scheme, which aims to provide fairness of allocation among users with heterogeneous demands, in a time-sharing system with multiple resource types, such as computing clouds [ghodsi2011dominant]. The services offered by these systems may include processing power, memory, and bandwidth, all of which are expected to be distributed fairly among the clients, according to some definition of fairness.
Although the problem has been well-studied for single resource types [keshav1997engineering, bertsekas1992data], prior to DRF, less attention was paid to the multi-resource scenarios. The general approach was to adopt well established solutions for single resource allocation, such as Max-min Fairness (MF) and its derivatives, for a fixed bundle with differing amounts of resources from each resource type; e.g. standardised virtual machines. But this approach leads to inefficiencies, since the needs of different users are heterogeneous, where one needs predominantly one resource type, the other needs another. In the end, users ended up with unused resources that come in the bundle with the resources they need, where at the same time there were other users in the system in demand of those unused resources.
@@
In the present paper, we analyse the DRF algorithm as introduced in [ghodsi2011dominant] in closer view, and based on the observations gathered thereby, introduce a new algorithm to approximate the DRF allocation in fewer steps. We call this new algorithm Precomputed Dominant Resource Fairness (PDRF), after its main working principle of precomputing the main allocation loop of DRF, and assigning resources to users without going through the tedious iterations of the loop.
One study that also needs to be mentioned at the onset, is [parkes2015beyond], the method and the findings of which closely resemble to that of the present study, with some key differences that allow us to implement our newly proposed algorithm. The analysis method and the algorithm they propose, the Extended Dominant Resource Fairness (EDRF) are similarly based on precomputing the final allocation of DRF, but they modeled the underlying linear programming problem, rather than the algorithm itself. They construct a theoretical framework that allows them to provide more rigorous proofs of DRF properties, which is the main objective of the article. We elaborate on this in Section 3.2.
@@ 2 Related Work
DRF is first proposed in [ghodsi2011dominant], where they introduce the new notion of establishing an MF allocation among dominant shares, rather than a fixed resource bundle of different resource types. They also provide an algorithm to carry out the distribution, and produce empirical data to demonstrate the performance of the algorithm. In their theoretical sections, they compare the virtues of the newly introduced allocation scheme to the ones existing in the economy literature, for this is as much a problem in the field of economy as it is in the field of computer science. Their results clearly demonstrate that DRF possesses various advantages over its alternatives.
The model is challenged in [dolev2012no]. They developed an alternative fairness notion, which they refer to as no justified complaints, based on the assumption that the fairness of distribution of resources should be assessed from a global system point of view, considering the bottleneck property of resources. A resource is a bottleneck resource, if the availability of this resource is limited and there is a race condition for accessing it. They developed the Bottleneck Based Fairness (BBF) algorithm to implement this paradigm as an alternative to DRF, with relative advantages of each algorithm in different use cases.
Not only challenged, but DRF has also been adopted, generalised, contextualised, or extended to different cases. For example, Gutman and Nisan [gutman2012fair], introduced an economical framework to compare and contrast the two systems described above. While doing so they also critically assessed both approaches from an economic point of view. Joe-Wong and colleagues [joe2013multiresource] investigates the trade-off between fairness and efficiency, providing a unifying framework with two families of fairness functions, DRF being a special case in one of them. Kash [kash2014no] takes a scenario in which the complete information on demands is not available, since users arrive and leave the system, and develops a Dynamic DRF (DDRF) to address the problem. The DDRF algorithm is improved for its performance in [li2015note].
Finally, DRF has been reformulated by Parkes and colleagues [parkes2015beyond] to EDRF, in order to study its properties more rigorously. EDRF is also a solution to a special case, where the resources, as well as user demands are well-divisible, i.e. fractions of demands can be allocated to users, but for obvious reasons this is not widely applicable to real life scenarios, where half a task does not provide any added utility to the users. In the same article, they also introduce Sequential Min-max Algorithm to address the indivisible demands scenario. One more contribution of EDRF is that it addresses the situations where there are users who submit zero demands for one or more resources.
@@ 3.1 Dominant Resource Fairness
The DRF allocation scheme is build on the concept of dominant shares. The dominant share for each user is defined as the maximum of the fraction of her demands for different resource types to their respective reserves in the system. As exemplified in [ghodsi2011dominant], in a system with $9$ CPUs and $18$ GB memory, the dominant share, $ds$ for user A, with the demand vector $\left\langle 1,4\right\rangle$ is memory, since the fractional demands, $fd$, of the user are $1/9$ and $4/18=2/9$, respectively. Similarly, the $ds$ of a user with the demand vector $\left\langle 3,1\right\rangle$ is for CPU, since $fd$’s are $3/9$ and $1/18$, respectively. The question then is, what is a fair allocation between the users, whose demands for the resources are defined as such.
If the number of tasks allocated to users A and B is denoted by $x$ and $y$, respectively, the DRF allocation is defined as the solution to the optimising problem below [ghodsi2011dominant, p. 4]:
@@
As seen in the example above, DRF adopts progressive filling strategy, by aiming to keep the growth of allocation among different agents at the same pace. DRF reduces to MF, and also to BBF [dolev2012no], in cases where all users’ dominant resources are the same.
@@ 3.2 Extended Dominant Resource Fairness
The study by [parkes2015beyond] introduces an alternative view of DRF. Instead of an iterative algorithm, they aim to precompute the resulting distribution, with a linear program. For doing so, they employ two notions in the setting. First, in a given user demand vector, the ratio between the entries for different resources is fixed. Second, the operating principles of progressive filling algorithm dictates that the pace at which the allocation for dominant shares of different users grow should be equated.
@@ 4 Cyclic Structure of the DRF Main Loop
In a system of $n$ users, the length of the cycle and the number of occurrences for each individual demand in the cycle are given by
@@
where $\lvert l\rvert$ denotes the length of the cycle, $\operatorname{lcm}DS$ denotes the least common multiple of the set of all dominant shares, $DS$, and $ds_{i}$ is the dominant share of user $i$. The individual occurrences of each demand within the cycle is given by the $i$-thth term of the summation for user $i$. In the cases where all other dominant shares are integer fractions of $ds^{*}$, like in the example above, since $\operatorname{lcm}DS$ is equal to $ds^{*}$, the latter part is equivalent to the ratio of each $ds_{i}$ to $ds^{*}$, and the length of the cycle is equal to the sum of all these ratios.
The ratio of lower demands to the $ds^{*}$ is still useful, in cases where integer fraction assumption does not hold. In those cases, further subcycles emerge, characterised by the single occurrence of $ds^{*}$, followed by all other demands occurring in the number of the ratio $ds_{i}$ to $ds^{*}$ (i.e. $ds^{*}/ds_{i}$), rounded up or down. The most frequently occurring subcycle is characterised by a single occurrence of the demand with $ds^{*}$, followed by all other demands occurring for times equal to $ds^{*}/ds_{i}$ rounded to the nearest integer. That is to say:
$\lvert l^{*}\rvert=\sum_{i=1}^{n}\left\lfloor\frac{ds^{*}}{ds_{i}}\right\rceil$
where $\lvert l^{*}\rvert$ denotes the length of, what we would like to call, the characteristic subcycle. The other subcycles are characterised by the occurrence of one or more of the lower dominant share demands equal to the ratio rounded in the opposite direction, the frequency of which is determined by the decimal part of the ratio of $ds^{*}$ to each dominant share, i.e. the $i$-thth term of the expression above for demand $i$. Trivially, the first subcycle will always round up, if the ratio has a decimal part.
Whenever a multiple of the decimal part reaches or exceeds an integer number, the ratio rounds up, otherwise to down. For example, in a system with two demands, whose dominant share ratio is $1.25$, the characteristic subcycle will consist of each demand taking $1$ allocation, and every $4th$ subcycle will include one more allocation for the lower dominant share, making it $1$ to $2$.
In cases where the decimal summations exceed an integer, such as in the case of a ratio of $1.4$, there will be $2$ different periods of rounding up; in this case, after the first subcycle, every $2nd$ and then $3rd$.
Ratios which have a decimal part greater than $0.5$, exceed integers more often than not, and thus will mostly round up, and round down in the fewer cases where they fail to.
Systems with more than two users, with all pairwise dominant share ratios having decimal parts, and not necessarily as nice as the ones presented in the examples above, lead to a cornucopia of subcycles, with differing and chaotically overlapping rounding periods for each demand. Luckily, we will not need to identify all of these for approximating DRF allocation.
@@ 5 Precomputed Dominant Resource Fairness
The pseudocode of PDRF is presented in Algorithm 1. A sample implementation can be accessed at this github repository [metin2025pdrf].
@@ 6 Results
$1000$ users are assigned random demands, and resource reserves are likewise assigned randomly. We tested for various numbers of resource types, demand intervals and resource reserve intervals. Here, we are presenting a sample which we consider representative. We run each test $1000$ times, and collected the averages, standard deviations, and maximums, presented in Tables 1 - 4. The scripts that we used to run the tests can be accessed at the same repository, here [metin2025pdrf].
@@ 9.2 Pareto Efficiency
The proof is based on the claim that ”progressive filling for DRF is equivalent to the algorithm presented in Figure 11111Since Figure $1$ is a scatter plot, we assume what is meant is Algorithm $1$. [ghodsi2011dominant, p. 13]. Although it is indicated in the preceding paragraph that the progressive filling algorithm eliminates the saturated demands, no removal condition is given in Algorithm $1$, and neither it is mentioned in the sections and examples describing the algorithm. This might be a minor point eluding the authors’ attention or it might have been assumed that the selection process automatically excludes those demands, but the resulting algorithm as such, cannot achieve Pareto Efficiency in, though rare might they be, certain demand distributions and resource reserves.
@@
An example we rather stumbled upon, which led to this analysis, consists of the same demand vectors in the original article, user A with the demand vector $\left\langle 1,4\right\rangle$ and user B with $\left\langle 3,1\right\rangle$, and the available capacities $\left\langle 59,19\right\rangle$. In such a setting, after 10 iterations, the algorithm halts allocating $\left\langle 2,8\right\rangle$ and $\left\langle 24,8\right\rangle$, respectively, and $\left\langle 33,3\right\rangle$ resources still available, which is sufficient for $3$ more allocations for user B, but since the least allocated dominant share belongs to A, and his demand is saturated, the algorithm cannot proceed. A short simulator script with this example can be found at the same github repository [metin2025pdrf].
@@ 9.3 Weighting Policies
A weighting policy, assigns weights to users, and allocates the resources in proportion to these weights. Fairness, in turn, is defined over the allocations scaled by these weights. In [ghodsi2011dominant], the question of weighting is addressed briefly, pointing to the fact that it suffices for such a generalisation to assign a weight vector to each user, $W_{i}=\left\langle w_{i1},w_{i2},w_{i3},\dots,w_{im}\right\rangle$, and compute the dominant shares with $ds_{i}=\max_{j}\left\{u_{ij}/w_{ij}\right\}$. In [parkes2015beyond], weights are further defined to be normalised, such that $\sum_{i}w_{ir}=1$, for each resource $r$, and they use another $\rho$” variable, at the construction of which the weights scale the normalised demands, to be used in their calculations. Again, it is notationwise involved and will not be elaborated on in more detail for the sake of simplicity.